# Active Preference Learning for Large Language Models

### v1

@@ 1 Introduction
Fine-tuning state-of-the-art LLMs using both of the aforementioned methods can require highly skilled domain experts, or expensive LLMs in the case of AI feedback, to produce the required preference data. In this work, we focus on how best to utilize the available preference labelling budget, specifically when using the DPO objective to avoid the need for RL. Instead of randomly selecting a large fixed number of prompts upfront and acquiring oracle labels for a subsequent fine-tuning process, we introduce an iterative data acquisition and fine-tuning loop that we refer to as Active Preference Optimization (APO). At each step, a batch of prompt/completion pairs is selected according to an acquisition function, oracle labels are acquired and then the model is improved with a cycle of fine-tuning. This loop is then repeated until some preference label budget is reached.
We develop a simple and effective acquisition function for prompt/completion pairs that uses the predictive entropy of the latest version of the model $p_{\theta_{t}}(y|x)$ and a measure of certainty of DPOs implicit preference model. Our active sampling approach biases the fine-tuning process towards correcting data points where the models implicit preference ranking is confidently wrong; leading to better learning outcomes. We also leverage an LLM oracle to provide preference labels online and use the latest version of the fine-tuned model to generate completions at each step.
In our experiments over multiple data sets using open source models with $\approx 1$ billion parameters, we demonstrate our approach improves win-rate performance of the fine-tuned model by on average  1-6%.
@@ 2 Direct Preference Optimization
$\mathcal{L}_{\phi}(\mathcal{X}_{P})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{X}_{P}}[\log\sigma(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l}))].$ (1)
@@
$\mathbb{E}_{x\sim\mathcal{X},y\sim p_{\theta}(y|x)}[r_{\phi}(x,y)]-\beta\mathcal{D}_{KL}(p_{\theta}(y|x)||p_{\theta_{0}}(y|x)).$ (2)
The second term here regularises the fine-tuned model using the KL-divergence to stay close to the state of the LLM before fine-tuning $p_{\theta_{0}}(y|x)$ The main rationale provided for this is to prevent the model from deviating too far from the distribution on which the reward model is accurate.
@@
$\mathcal{L}_{\theta}(\mathcal{X}_{P})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{X}_{P}}\left[\log\sigma\left(\hat{r}(x,y_{w})-\hat{r}(x,y_{l})\right)\right],$ (4)
@@
$\hat{r}(x,y)=\beta\frac{p_{\theta}(y|x)}{p_{\theta_{0}}(y|x)}.$ (5)
This formulation has the distinct advantage of not requiring the explicit reward modeling step and avoids the need to perform any reinforcement learning. Furthermore it has been shown to outperform RLHF across a range of experiments (Rafailov et al., 2023).
@@ 3 Active Preference Learning
We first outline our general active learning training procedure before introducing our acquisition functions for data selection. Informally, active learning is a paradigm in machine learning that aims to iteratively select the most useful datapoints during training using the current state of the model. Specifically we are interested in the setting of pool-based active learning which involves selecting a subset of observations from a closed pool of unlabeled data (Ren et al., 2021).
Our APO training algorithm consists of iterations of the following scheme: randomly sample a large batch of prompts; generate pairs of completions for each prompt according to the latest version of the fine-tuned $p_{\theta_{t}}(y|x)$; rank the prompt/completion pairs according to our acquisition function; select the highest ranking subset as a batch of preference pairs for fine-tuning; query the oracle to get preference labels on this batch and, finally, fine-tune $p_{\theta_{t}}(y|x)$ using the preference labels to produce $\theta_{t+1}$. This process is repeated until some preference labelling budget has been reached.
@@
Algorithm 1 Active Fine-tuning Procedure
@@ 3.1 Acquisition functions
In selecting scoring methods (step 8 in 1) we aim for options that are straightforward to implement and do not require modifications to the model architectures or the fine-tuning procedure itself. This allows for a drop in addition to existing implementations. As a result, we propose using the predictive entropy of $p_{\theta_{t}}(y|x)$ in combination with a measure of certainty under the Bradley-Terry preference model, which leverages the implicit reward model in DPO.
@@ 3.1.1 Entropy of the language model
$\approx\frac{1}{N}\sum_{n=1}^{N}\log p_{\theta}(y^{n}|x),$ (8)
@@ 3.1.2 Preference model certainty
We prioritize prompt/completion pairs with higher differences during the selection of data points for fine-tuning. The hypothesis here is that data points with high values provide valuable learning opportunities. Should the model’s implicit preference predictions diverge from the oracle’s evaluation, especially with high certainty, prioritising these discrepancies when fine-tuning can enhance model performance.
This is a natural choice given the behaviour of the DPO training objective (equation 4). Consider the gradient update with respect to the parameters $\theta$
$\small\nabla_{\theta}\mathcal{L}_{\theta}=-\beta\mathbb{E}_{\mathcal{X}_{P}}\left[w(\nabla_{\theta}\log p_{\theta}(y_{w}|x)-\nabla_{\theta}\log p_{\theta}(y_{l}|x))\right],$ (10)
where $w=\sigma(\hat{r}(x,y_{l})-\hat{r}(x,y_{w}))$ weights each sample $(x,y_{w},y_{l})\sim\mathcal{X}_{P}$. This gradient update can be interpreted as weighting examples by how incorrectly the implicit reward model is, while accounting for the strength of the KL constraint. Early in fine-tuning, when the implicit preference model is still likely to be wrong often, our proposed acquisition strategy prioritises examples that result in substantial gradient updates, which we find to accelerate learning progress and final performance in our experiments in section 5.
@@ 3.1.3 A hybrid approach
In practise we can combine both entropy and preference certainty as complimentary metrics for scoring data to exploit the strengths of both. The intuition is that higher entropy prompts are more likely to give incorrect predictions from the implicit preference model. In our experiments we first select a relatively large batch of prompts and rank them by the entropy. We then take the top subset of prompts ranked by entropy and generate the required completion pairs before scoring and ranking according to preference certainty. Finally, we take the top subset of prompt/completion pairs ranked by preference certainty and add them to our preference dataset for fine-tuning.
@@ 3.2 Choice of oracle
In algorithm 1, we require an oracle to provide preference judgments on pairs of completions for fine-tuning (step 10) and for evaluating against a held-out test dataset (step 13). Since we aim to generate completions using the latest version of the model at each data acquisition step, using pre-labeled datasets is not feasible. Additionally, relying on human judgments is impractical due to the need for multiple experiments with different datasets, models, acquisition functions, and seeds. To address this, we turn to state-of-the-art closed-source models offered by OpenAI. The question then becomes whether these models are suitable and, if so, which model should be chosen and how should it be prompted?
@@ 3.2.2 Choice of base model
A downside of using GPT-4 as our oracle model is the cost and high latency. A far more economical choice would be to use older versions of models such as GPT-3.5. We ran an analysis where we generated preference labels twice for both GPT-3.5 and GPT-4 on a set of 50 prompts and completions sampled sampled from the fine-tuning from human preferences dataset (Ziegler et al., 2019). Unfortunately we found that GPT-4 was far more consistent ($>$90%) than GPT-3.5-turbo (only $\sim$60%) at a range of sampling temperatures - see figure 1. We therefore chose to use GPT-4111Specifically model version gpt-4-1106-preview as the oracle for our experiments and adjusted our budget of evaluations appropriately. To note, our analysis assumes the same prompt for both models; we leave to future work further prompt engineering to improve the evaluation quality and consistency of smaller, more economical models.
@@ 3.3 Fine-tuning details
Here we discuss in more detail the implementation details for the fine-tuning step (12) in algorithm 1. We adopt the most straight-forward implementation, which is to re-initialise $\theta_{t}$ to $\theta_{0}$ at each time step $t$ and fine-tune to convergence, sampling uniformly from all previously acquired preference data $\mathcal{X}_{p}$. This is consistent with previous work on deep active learning (Gal et al., 2017) and relies on the assumption that cost (in time and/or money) of acquiring oracle labels outweighs the cost of fine-tuning again on all acquired data after each new batch of labels is acquired. The focus of our experiments is to isolate the differences in performance caused by the different acquisition vs randomly acquiring data.
We also must set the number of fine-tuning epochs to do at each fine-tuning step $t$. We base this choice on an empirical analysis of the number of epochs it took on average for our choice of models to converge at different dataset sizes. Convergence was measured on the performance against a validation dataset. We analysed loss and win-rate curves for the different model and dataset combinations - see appendix C.1 for details.
@@ 4 Related Work
Another orthogonal application of active learning in the setting of improving pre-trained langauge model performance is the active sampling of few shot examples for prompt stuffing (Margatina et al., 2023). In this work, the authors use acquisition functions based on different uncertainty, diversity and similarity scores of the language model across datasets of few shot examples to determine which examples are best to reference in the prompt to improve performance. Although similar in spirit to our work, they don’t consider actually updating the parameters of the model using preference labelled data.
An alternative approach to active learning is that of data pruning. In (Marion et al., 2023), pruning heuristics are applied to filter the data used in the first stage of unsupervised LLM pre-training. This leads to improved performance on downstream tasks versus the LLMs pre-trained on the full dataset. Over $50\%$ of the data can be pruned while still leading to improvements. This work does not directly consider the impact of such pruning techniques for the preference fine-tuning stage, but some of their perplexity based heuristics could represent viable alternatives to our acquisition strategies.
@@ 5 Experiments
The focus of our experiments is to determine if more active sampling during the fine-tuning process can bring us gains in data efficiency when dealing with limited labelling budgets; in terms of the rate of learning and the final performance achieved. We compare three different acquisition functions: random, certainty and entropy + certainty (as discussed in section 3.1). We evaluate across two different open source large language models and two different datasets used in recent related work. We also want to gather some qualitative findings about the characteristics of the datapoints being acquired under the different schemes.
@@ 5.1 Datasets
In line with recent work (Ziegler et al., 2019; Rafailov et al., 2023) we focus on two distinct datasets for our experiments; IMDB and TLDR. IMDB is a dataset of movie reviews where the task is to complete a positive review given the start of a review. TLDR, a more difficult task, is a dataset of Reddit posts where the task is to provide a summary of the post. Table 1 provides a summary of the dataset details. TLDR also provides human provided completions that can be used for evaluation.
We provide further details on dataset pre-processing in appendix B.
@@
Task
Complete reviews according to preference
Generate summaries according to preferences
@@
ADAM lr: 1e-06
ADAM lr: 1e-06
as per [v1] arxiv version of DPO paper (Rafailov et al., 2023)
@@ 5.2 Models
For both IMDB and TLDR we used large transformer based architectures as introduced in section. See table 1 for a summary of the models and main hyper-parameters used in both cases. For IMDB, the GPT-2 base transformer model provided by Hugging Face [^2] Downloaded pre-trained base model from https://huggingface.co/edbeeching/gpt2-large-imdb was pre-trained on the WebText corpus (Radford et al., 2019) and has 12 layers with 768 dimensions, with 12 attention heads. It was also further trained in an unsupervised way on the full IMDB dataset.
For TLDR we use the Pythia  [^3] Pre-trained base model from https://huggingface.co/pvduy/pythia-1B-sft-summarize-tldr class of transformer model (Biderman et al., 2023) that has an architecture similar to GPT-3, with 805M parameters, 16 layers with 2048 dimensions and 8 attention heads. We ran our fine-tuning on single 40GB RAM A100 and 48GB 600 ADAs GPUs throughout our experiments.
@@ 5.4 Acquisition sampling
In our experiments we first randomly sample $S=4000$ for IMDB and $S=2048$ for TLDR for our preference certainty acquisition runs. When doing entropy + preference certainty, we first randomly sample $J\times S$ prompts, rank them by entropy and take the top $S$ prompts to generate completions before further scoring and ranking by preference certainty. We use $J=8$ for IMDB and $J=4$ for TLDR. We use $N=8$ samples when approximating the entropy. For all experiments we set the final acquisition batch size to $M=128$.
@@ 5.5 Evaluation
The choice of oracle for providing labels and evaluating on the test data is GPT-4. Details of the prompts are provided in appendix A. Our prompts specify a task specific preference, but also consider grammatical correctness and consistency. Our evaluation approach on held out test prompts uses head-to-head win-rate comparisons versus completions sampled from the pre-trained model from the start of training $p_{\theta_{0}}(y|x)$ for IMDB. In the case of TLDR, we replace the pre-trained model completions with the human-provided completions that were available on the hold out test data. Due to the significant cost associated to using GPT-4 as the oracle for evaluation, we don’t evaluate after every single data acquisition step. Each evaluation is done against 1024 test prompts.

### v2

@@ 1 Introduction
Fine-tuning state-of-the-art LLMs using both of the aforementioned methods can require highly skilled domain experts, or expensive LLMs in the case of AI feedback, to produce the required preference data. In this work, we focus on how best to utilize the available preference labelling budget, specifically when using the DPO objective to avoid the need for RL. Instead of randomly selecting a large fixed number of prompts upfront and acquiring oracle labels for a subsequent fine-tuning process, we introduce an iterative data acquisition and fine-tuning loop that we refer to as Active Preference Learning (APL). At each step, a batch of prompt/completion pairs is selected according to an acquisition function, oracle labels are acquired and then the model is improved with a cycle of fine-tuning. This loop is then repeated until some preference label budget is reached.
We develop a simple and effective acquisition function for prompt/completion pairs that uses the predictive entropy of the latest version of the model $p_{\theta_{t}}(y|x)$ and a measure of certainty of DPO’s implicit preference model. Our active sampling approach biases the fine-tuning process towards correcting data points where the models implicit preference ranking is confidently wrong; leading to better learning outcomes. We also leverage an LLM oracle to provide preference labels online and use the latest version of the fine-tuned model to generate completions at each step.
In our experiments over multiple data sets using open source models with $\approx 1$ billion parameters, we demonstrate our approach improves the win-rate performance of the fine-tuned model by on average  1-6%.
@@ 2 Direct Preference Optimization
$\mathcal{L}_{\phi}(\mathcal{X}_{P})=-\mathbb{E}_{\mathcal{X}_{P}}[\log\sigma(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l}))].$ (1)
@@
$\mathbb{E}_{x\sim\mathcal{X},y\sim p_{\theta}(y|x)}[r_{\phi}(x,y)]-\beta\mathrm{KL}(p_{\theta}(y|x)||p_{\theta_{0}}(y|x)).$ (2)
The second term here regularises the fine-tuned model using the KL-divergence to stay close to the state of the LLM before fine-tuning $p_{\theta_{0}}(y|x)$. The main rationale provided for this is to prevent the model from deviating too far from the distribution on which the reward model is accurate.
@@
$\mathcal{L}_{\theta}(\mathcal{X}_{P})=-\mathbb{E}_{\mathcal{X}_{P}}\left[\log\sigma\left(\hat{r}(x,y_{w})-\hat{r}(x,y_{l})\right)\right],$ (4)
@@
$\hat{r}(x,y)=\beta\log\frac{p_{\theta}(y|x)}{p_{\theta_{0}}(y|x)}.$ (5)
This formulation has the distinct advantage of not requiring the explicit reward modeling step and avoids the need to perform any reinforcement learning. Furthermore, it has been shown to outperform RLHF across a range of experiments (Rafailov et al., 2023).
@@ 3 Active Preference Learning
We first outline our active learning training procedure before introducing our acquisition functions for data selection. Informally, active learning is a paradigm in machine learning that aims to iteratively select the most useful datapoints during training using the current state of the model. Specifically, we are interested in the setting of pool-based active learning which involves selecting a subset of observations from a closed pool of unlabeled data (Ren et al., 2021).
Our APL training algorithm consists of iterations of the following scheme: randomly sample a large batch of prompts; generate pairs of completions for each prompt according to the latest version of the fine-tuned $p_{\theta_{t}}(y|x)$; rank the prompt/completion pairs according to our acquisition function; select the highest ranking subset as a batch of preference pairs for fine-tuning; query the oracle to get preference labels on this batch and, finally, fine-tune $p_{\theta_{t}}(y|x)$ using the preference labels to produce $\theta_{t+1}$. This process is repeated until some preference labelling budget has been reached.
@@
Algorithm 1 Active Preference Learning Procedure
@@ 3.1 Acquisition functions
In selecting scoring methods (step 8 in 1) we aim for options that are straightforward to implement and do not require modifications to the model architectures or the fine-tuning procedure itself. This allows for a drop in addition to existing implementations. As a result, we propose using the predictive entropy of $p_{\theta_{t}}(y|x)$ as well as a measure of certainty under the Bradley-Terry preference model, which leverages the implicit reward model in DPO.
@@ 3.1.1 Entropy of the language model
$\approx-\frac{1}{N}\sum_{n=1}^{N}\log p_{\theta}(y^{n}|x),$ (8)
@@ 3.1.2 Preference model certainty
We prioritize prompt/completion pairs with higher differences during the selection of data points for fine-tuning. Our hypothesis is that data points with high values provide valuable learning opportunities. Should the model’s implicit preference predictions diverge from the oracle’s evaluation, especially with high certainty, prioritising these discrepancies when fine-tuning can enhance model performance.
This choice is well motivated by the behaviour of the DPO training objective (equation 4). Consider the gradient update with respect to the parameters $\theta$
$\nabla_{\theta}\mathcal{L}_{\theta}=-\beta\mathbb{E}_{\mathcal{X}_{P}}\left[w(\nabla_{\theta}\log p_{\theta}(y_{w}|x){-}\nabla_{\theta}\log p_{\theta}(y_{l}|x))\right],$ (10)
where $w=\sigma(\hat{r}(x,y_{l})-\hat{r}(x,y_{w}))$ weights each sample $(x,y_{w},y_{l})\sim\mathcal{X}_{P}$. This gradient update can be interpreted as weighting examples by how incorrectly the implicit reward model is while accounting for the strength of the KL constraint. Early in fine-tuning, when the implicit preference model is still likely to be wrong often, our proposed acquisition strategy prioritises examples that result in substantial gradient updates, which we find to accelerate learning progress and lead to an improvement in the final performance in our experiments in section 5.
@@ 3.1.3 A hybrid approach
In practise we can combine both entropy and preference certainty as complimentary metrics for scoring data to exploit the strengths of both. Our hypothesis is that higher entropy prompts are more likely to give incorrect predictions from the implicit preference model. In our experiments for this hybrid approach, we first select a relatively large batch of prompts and rank them by the entropy. We then take the top subset of prompts ranked by entropy and generate the required completion pairs before scoring and ranking according to preference certainty. Finally, we take the top subset of prompt/completion pairs ranked by preference certainty and add them to our preference dataset for fine-tuning.
@@ 3.2 Choice of oracle
Algorithm 1 requires an oracle to provide preference judgments on pairs of completions for fine-tuning (step 10) and for evaluating against a held-out test dataset (step 13). Since we aim to generate completions using the latest version of the model at each data acquisition step, using pre-labeled datasets is not feasible. Additionally, relying on human judgments is impractical due to the need for multiple experiments with different datasets, models, acquisition functions, and seeds. To address this, we turn to state-of-the-art closed-source models offered by OpenAI. The question then becomes whether these models are suitable and, if so, which model should be chosen and how should it be prompted?
@@ 3.2.2 Choice of base model
A downside of using GPT-4 as our oracle model is the cost and high latency. A far more economical choice would be to use older versions of models such as GPT-3.5. We ran a simple analysis where we generated preference labels twice for both GPT-3.5 and GPT-4 on a set of 50 prompts and completions sampled from the fine-tuning from human preferences dataset (Ziegler et al., 2019). Unfortunately, we found that GPT-4 was far more consistent ($>$90%) than GPT-3.5-turbo (only $\sim$60%) at a range of sampling temperatures - see figure 1. We, therefore, chose to use GPT-4111Specifically model version gpt-4-1106-preview as the oracle for our experiments and adjusted our budget of evaluations appropriately. To note, our analysis assumes the same prompt for both models; we leave to future work to further prompt engineering to improve the evaluation quality and consistency of smaller, more economical models.
@@ 3.3 Fine-tuning details
Here we discuss in more detail the implementation details for the fine-tuning step (12) in algorithm 1. We adopt the most straight-forward implementation, which is to re-initialise $\theta_{t}$ to $\theta_{0}$ at each time step $t$ and fine-tune to convergence, sampling uniformly from all previously acquired preference data $\mathcal{X}_{p}$. This is consistent with previous work on deep active learning (Gal et al., 2017) and relies on the assumption that the cost (in time and/or money) of acquiring oracle labels outweighs the cost of fine-tuning again on all acquired data after each new batch of labels is acquired. The focus of our main experiments in section 5 is to isolate the differences in performance caused by the different acquisition vs randomly acquiring data. In Appendix D, we discuss adapting our approach for online learning and present some provisional results.
We must also set the number of fine-tuning epochs to perform at each fine-tuning step $t$. We base this choice on an empirical analysis of the number of epochs it took on average for our choice of models to converge at different dataset sizes. Convergence was measured on the performance against a validation dataset. We analysed loss and win-rate curves for the different model and dataset combinations - see Appendix E for details.
@@ 4 Related Work
Another orthogonal application of active learning in the setting of improving pre-trained LLM performance is the active sampling of few shot examples for prompt stuffing (Margatina et al., 2023). In this work, the authors use acquisition functions based on different uncertainty, diversity and similarity scores of the language model across datasets of few-shot examples to determine which examples are best to reference in the prompt to improve performance. Although similar in spirit to our work, they don’t consider updating the parameters of the model using preference-labelled data.
An alternative active learning approach is data pruning. In (Marion et al., 2023), pruning heuristics are applied to filter the data used in the first stage of unsupervised LLM pre-training. This leads to improved performance on downstream tasks versus the LLMs pre-trained on the full dataset. Over $50\%$ of the data can be pruned while still leading to improvements. This work does not directly consider the impact of such pruning techniques for the preference fine-tuning stage, but some of their perplexity based heuristics could represent viable alternatives to our acquisition strategies.
Finally, a research theme adjacent to active learning that can also reduce the amount of preference labels required is that of self-play fine-tuning (Chen et al., 2024; Yuan et al., 2024). These works focus on how to bootstrap $p_{\theta_{t}}(y|x)$ during fine-tuning to provide preference labels, or to act as a reward model, as opposed to trying to make better use of oracle resources. This in principle could be combined with our active preference learning approach and so we consider it complimentary.
@@ 5 Experiments
The focus of our experiments is to determine if more active sampling during the fine-tuning process can bring us gains in data efficiency when dealing with limited labelling budgets; in terms of the rate of learning and the final performance achieved. We compare four different acquisition configurations: random, entropy, certainty and entropy + certainty (as discussed in section 3.1). We evaluate across two different open source large language models and two different datasets used in recent related work. We also gather some qualitative findings about the characteristics of the datapoints being acquired under the different schemes, which we discuss further in 5.6.1.
@@ 5.1 Datasets
In line with recent work (Ziegler et al., 2019; Rafailov et al., 2023) we focus on two distinct datasets for our experiments; IMDB and TLDR. IMDB is a dataset of movie reviews where the task is to complete a positive review given the start of a review. TLDR, a more difficult task, is a dataset of Reddit posts where the task is to provide a summary of the post. Table 1 provides a summary of the dataset details. TLDR also provides human-provided completions that can be used for evaluation. We provide further details on dataset pre-processing in Appendix B.
@@
Task Complete reviews according to preference Generate summaries according to preferences
@@ 5.2 Models
For both IMDB and TLDR we use relatively large transformer based architectures. See table 1 for a summary of the models and main hyper-parameters used in both cases. For IMDB, the GPT-2 base transformer model provided by Hugging Face [^2] Downloaded pre-trained base model from https://huggingface.co/edbeeching/gpt2-large-imdb was pre-trained on the WebText corpus (Radford et al., 2019) and has 12 layers with 768 dimensions, with 12 attention heads. It was also further trained in an unsupervised way on the full IMDB dataset. For TLDR, we use the Pythia [^3] Pre-trained base model from https://huggingface.co/pvduy/pythia-1B-sft-summarize-tldr class of transformer model (Biderman et al., 2023) that has an architecture similar to GPT-3, with 805M parameters, 16 layers with 2048 dimensions and 8 attention heads. We ran our fine-tuning on single 40GB RAM A100 and 48GB 600 ADAs GPUs throughout our experiments.
@@ 5.4 Acquisition sampling
In our experiments we first randomly sample $S=4000$ for IMDB and $S=2048$ for TLDR for our entropy only and preference certainty only acquisition runs. When doing entropy + preference certainty, we first randomly sample $J\times S$ prompts, rank them by entropy and take the top $S$ prompts to generate completions before further scoring and ranking by preference certainty. We use $J=8$ for IMDB and $J=4$ for TLDR. We use $N=8$ samples when approximating the entropy. For all experiments we set the final acquisition batch size to $M=128$.
@@ 5.5 Evaluation
We use GPT-4 as the oracle for providing labels and evaluating the test data. Details of the prompts are provided in Appendix A. Our prompts specify a task-specific preference but also consider grammatical correctness and consistency. Our evaluation approach on held-out test prompts uses head-to-head win-rate comparisons versus completions sampled from the pre-trained model from the start of training $p_{\theta_{0}}(y|x)$ for IMDB. For TLDR, we replaced the pre-trained model completions with the human-provided completions that were available on the hold-out test data. Due to the significant cost of using GPT-4 as the oracle for evaluation, we don’t evaluate after every single data acquisition step. Each evaluation is done against 1024 test prompts.